One pump fills a tank 2 times as fast as another pump. If the pumps work together, they fillthe tank in 8 minutes. How long does it take for each pump to fill the tank? It takes the faster pump Γ minutes to fill the tank when working alone and it takes the slower pump minutes
Answer #1

Let the faster pump can fill the tank in x minute.

The slower pump can fill the tank in 2x minute.

(x * 2x)/(x + 2x) = 8

or, 2x^2/3x = 8

or, 2x^2 = 24x

or, 2x^2 - 24x = 0

or, 2x(x - 12) = 0

2x = 0 or, x - 12 = 0

or, x = 0 or, x = 12

So the faster pump can fill the tank in 12 minute.

The slower pump can fill the tank in 2 * 12 = 24 minute.